OpenStudy (anonymous):

0.095 must be multiplied by what power of ten to make a whole number

OpenStudy (anonymous):

well: 0.095 x 10^1 = 0.95 <--- not a whole number yet 0.095 x 10^2 = 9.5 <--- still not a completely whole number 0.095 x 10^3 = 95 <--- bingo! So in order to get 0.095 to become a whole number, you need to multiply it by a 10 to the third power or 10^3